OpenStudy (anonymous):

2. What is the solution of the system? 7x + 5y = 19 –7x – 2y = –16 (1 point) (2, 1) (1, 2) (7, 5) (–7, –2)

OpenStudy (natalieee):

just plug it in where the x and y's are to get your answer.

OpenStudy (natalieee):

2,1

OpenStudy (natalieee):

medal me

OpenStudy (anonymous):

thx

OpenStudy (natalieee):

your welcome (:

OpenStudy (anonymous):

addition/elimination method 7x + 5y = 19 - 7x - 2y = - 16 ----------------add 3y = 3 y = 1 Now substitute 1 in for y in either of the two equations... 7x + 5y = 19 7x + 5(1) = 19 7x + 5 = 19 7x = 19 - 5 7x = 14 x = 2 The answer is (2,1)